The value of ΔG° at 241.0°C for the formation of phosphorous trichloride from its constituent elements,
P2(g) + 3Cl2(g) → 2PCl3(g)
is ________ kJ/mol. At 25.0°C for this reaction, ΔH° is -720.5 kJ/mol, ΔG° is -642.9 kJ/mol, and ΔS° is -263.7 J/K.
Answer #1

ΔHo = -720.5 KJ/mol

ΔSo = -263.7 J/mol.K

= -0.2637 KJ/mol.K

T= 241.0 oC

= (241.0+273) K

= 514 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -720.5 - 514.0 * -0.2637

ΔGo = -585.0 KJ/mol

Answer: -585.0 KJ/mol